# LLMs in the Heart of Differential Testing: A Case Study on a Medical Rule Engine

### v1

@@ 6.1. Hallucination Effect
(1) Semantic alterations.This limitation refers to small semantic changes such as transforming “violation_case” to “violating_case”. While seemingly minor, such a change is considered incorrect.
(2) Different levels of detail.We observe discrepancies in the level of detail between the expected and generated tests. This manifests in three common scenarios:
(a) Missing test types: In some instances, the LLMs fail to generate all specified test types. For example, out of the three expected test types (Pass, Fail, and NotApplied), only two might be generated (Pass and Fail).
(b) Additional tests: The LLMs might produce redundant tests, unnecessarily generating variations of the same scenario (e.g., generating “invalid_case”, “invalid_case_1”, and “invalid_case_2”).
@@
6.2. No LLM Subsumes Another
Our differential testing results reveals that the most effective LLM (i.e., GPT-3.5) does not uncover the most mismatches.
@@ 6.2. No LLM Subsumes Another
We identify three specific cases:
(1) inconsistency in handling rule versions,
(2) inconsistency in handling date formats,and
(3) inconsistency in handling variable dependencies.
@@
For instance, while Mistral successfully detects all rules involving the “ds” variable, it overlooks cases triggering a 500 status code—unlike Mixtral.
@@
While any of the LLMs would be suited for detecting mismatches, the optimal choice for a developer or researcher requires deeper understanding of the specific inconsistency types each model excels at identifying and the concrete usage scenario.
Additionally, it is important to remain open to exploring other LLMs that potentially improve on detecting specific inconsistencies or offer novel insights.
6.3. Using LLMs Instead of Other Test Generation Techniques
@@
While such solvers are efficient in finding satisfying and violating variable assignments, they may not capture the “meaning” of variable names and relationships among variables and constraints. Nonetheless, a dedicated empirical evaluation is needed to compare LLMs with other constraint-solving approaches (e.g., based on search algorithms and satisfiability modulo theories (SMT) solvers).
6.4. LLMs Might Not Find Simple Faults
@@
6.5. Mismatches Faults
In RQ3, we discussed the differential testing results in terms of rule result matches and mismatches.
Our study revealed that LLMeDiff serves as an effective technique for uncovering disparities within the rule engines.
However, it is important to emphasize that LLMeDiff’s task is limited to flagging these mismatches.
There is a potential for false positives and negatives (e.g., data overlaps between rule versions—erroneously leading to Pass), which complicates the interpretation of the results.
False positives might stem from the reference implementation (Dvare++) not accurately reflecting the desired behavior of GURI.
Therefore, determining whether these differences signify real faults in the rule engine requires manual root cause analysis by the CaReSS developers.
While such false positives add work for the developers, we assume that checking whether a mismatch is a false positive will neither require much time nor effort, as a domain expert can quickly identify a false positive.
@@ 6.5. Mismatches Faults
6.6. Improvement on the State of the Practice
@@ 6.7. Generality of the Results
(1) Our findings are relevant to other health registries in Norway (e.g., rthe Norwegian Patient Registry) and similar real-world contexts, such as the Norwegian toll and customs, which implement similar rule-based systems.
@@
7. Related Work
A typical software testing challenge in many contexts is the absence of precise and reliable test oracles [8], i.e., checking whether a test case’s outcome meets the expected outcome using a test oracle. In many real-world situations (including our context of testing medical rule engines), such an oracle either doesn’t exist or is impractical due to resource constraints. Metamorphic and differential testing techniques have been employed in the literature to deal with the test oracle problem [35, 15]. In this paper, we applied a typical differential testing setup to support testing GURI, i.e., using Dvare++ as a simplified reference implementation of GURI. The key idea is to use GURI and Dvare++ to check whether both systems produce consistent outputs when provided identical inputs generated by LLMs. Such a comparison allows for identifying inconsistencies, hence identifying possible issues in the implementation of GURI. As a result, we address the test oracle problem to some extent.
Recently, LLMs have become popular in software testing [43], e.g., focusing on unit testing [16, 47], system input generation [12, 11], property-based test generation [42], and test oracle generation [40]. Our work is similar to these works, focusing on generating tests with LLMs that let medical rules yield Pass, Fail, or NotApplied results; however, the application context is specific to the medical domain.
Consequently, our work is related to the works that aim to explore the potential of LLMs in the medical domain. For example,
Zhang et al. [49] evaluates the performance of LLMs, including GPT-3.5, GPT-4, Falcon, and LLama 2, in the classification of medical terms, specifically using a metastatic cancer dataset to assess reasoning consistency. Wornow et al. [46] evaluates models such as GPT-4 on clinical text and concludes that many models are trained on small or broad biomedical datasets, emphasizing the importance of a better evaluation system that aligns with healthcare standards. In general, surveys [34, 38, 10] provide insights into the potential and challenges of LLMs in healthcare applications and highlight that the practical use of LLMs remains limited. Our work with the CRN is a step towards the practical deployment of LLMs for testing their software systems.
In the past, we developed a model-based framework to model the medical rules for GURI [44] with the Unified Modeling Language (UML) and Object Constraint Language (OCL). Next, we developed an impact analysis approach [45] to study the impact of rule changes in response to GURI’s evolution (e.g., changes in regulations). In addition, we developed a search-based approach to automatically refactor rules [30] to enhance rule understandability and maintainability. However, these approaches do not help finding problems in the modeled rules. More recently, four studies advanced the testing at the CRN, mainly focusing on the GURI subsystem [26, 27, 21, 29].
@@ 7. Related Work
Laaber et al. [27] address the practical complexities of testing an evolving system, offering insights applicable to healthcare registries. Moreover, a case study focused on applying EvoMaster [3] (an AI-based testing tool) to assess its effectiveness regarding code coverage, error detection, and domain-specific rule coverage at the CRN [26].
Meanwhile, we also proposed EvoClass [21] by extending EvoMaster with a ML classifier, showing significant cost reductions without compromising effectiveness. Lastly, EvoCLINICAL [29], a cyber-cyber digital twin for GURI, was proposed, which with fine-tuning through active transfer learning, can deal with system evolution. In contrast to these existing works, we make new contributions by addressing the test oracle problem associated with rule execution results by employing differential testing. Moreover, we tackle the challenge of generating electronic health records (EHRs) by leveraging the capabilities of LLMs.
8. Conclusions
@@
Acknowledgements.
The research presented in this paper has received funding from The Research Council of Norway (RCN) under the project 309642 and has benefited from the Experimental Infrastructure for Exploration of Exascale Computing (eX3), which is supported by the RCN project 270053.
The Norwegian Ministry of Education and Research supports Erblin Isaku’s PhD.

### v4

@@ VI-A Hallucination Effect
(1) Semantic alterations.This limitation refers to small semantic changes, such as transforming “violation_case” to “violating_case”, that lead to incorrect outputs.
(2) Different levels of detail.We observe three discrepancies between the expected and generated tests:
(a) Missing test types: In some instances, the LLMs only generate a subset of the three specified test types (i.e., Pass, Fail, and NotApplied).
(b) Additional tests: The LLMs produce redundant tests, unnecessarily generating variations of the same scenario (e.g., generating “invalid_case”, “invalid_case_1”, and “invalid_case_2”).
@@
VI-B No LLM Subsumes Another
Our differential testing results reveal that the most effective LLM (i.e., GPT-3.5) does not uncover the most mismatches.
@@ VI-B No LLM Subsumes Another
We identify three specific cases leading to inconsistencies when handling
(1) rule versions,
(2) date formats,and
(3) variable dependencies.
@@
For instance, while Mistral successfully detects all rules involving the “ds” variable, it misses cases triggering a 500 status code—unlike Mixtral.
@@
While any of the LLMs would be suited for detecting mismatches, the optimal choice requires a deeper understanding of the specific inconsistency types each model excels at identifying and the concrete usage scenario.
Additionally, exploring other LLMs might improve detecting specific inconsistencies or offer novel insights.
VI-C Using LLMs Instead of Other Test Generation Techniques
@@
While such solvers are efficient in finding satisfying and violating variable assignments, they may not capture the “meaning” of variable names and relationships among variables and constraints.
This capability is not limited to assigning appropriate values to variables.
It also ensures that the generated test inputs consider the logical and domain-specific constraints between variables, such as temporal dependencies (e.g., diagnosis date preceding treatment date) or valid combinations of attributes (e.g., certain diagnoses requiring specific procedures).
Nonetheless, a dedicated empirical evaluation is needed to compare LLMs with other constraint-solving approaches (e.g., based on search algorithms and satisfiability modulo theories (SMT) solvers).
VI-D LLMs Might Not Find Simple Faults
@@
VI-E Mismatches Faults
In RQ3, we analyzed differential testing results and demonstrated that LLMeDiff effectively uncovers disparities within rule engines.
However, its role is limited to flagging mismatches, which may include false positives and negatives (e.g., data overlaps between rule versions leading to Pass).
Mismatches are not inherently good or bad; their value lies in their ability to provide actionable insights.
Useful mismatches highlight faults in rule logic or system behavior, while non-actionable mismatches, such as those caused by invalid input formats, indicate areas for improving input validation or test generation.
While more mismatches may suggest better test coverage, they also increase the workload for manual analysis, requiring developers to discern actionable issues.
Thus, the focus should not be on maximizing or minimizing mismatches but on generating meaningful tests that reveal actionable insights and improve system reliability.
@@ VI-E Mismatches Faults
VI-F Improvement on the State of the Practice
@@ VI-G Generality of the Results
(1) Our findings are relevant to other health registries in Norway (e.g., the Norwegian Patient Registry) and similar real-world contexts, such as the Norwegian toll and customs, which implement similar rule-based systems.
@@
VII Related Work
A prominent software testing challenge is the absence of precise and reliable test oracles [7], i.e., checking a test case’s actual outcome with the expected outcome using a test oracle. In many real-world situations (including our context), such an oracle either doesn’t exist or is impractical due to resource constraints.
Metamorphic and differential testing techniques are solutions to the test oracle problem [35, 18].
In this paper, we applied differential testing to GURI using Dvare++ as a simple reference implementation.
The idea is to check whether both systems produce consistent outputs on identical inputs generated by LLMs and identify inconsistencies stemming from potential implementation issues of GURI.
As a result, we address the test oracle problem to some extent.
Recently, LLMs have gained popularity in software testing [42], including unit testing [19, 46], system input generation [11, 10], property-based test generation [41], and test oracle generation [39]. Our work focuses on generating tests with LLMs to produce Pass, Fail, or NotApplied results, specifically for the medical domain. Related work includes Zhang et al. [48], who evaluate LLMs, such as GPT-4 and Llama 2, for medical term classification using metastatic cancer data, and Wornow et al. [45], who assess models like GPT-4 on clinical text, emphasizing the need for better evaluation aligned with healthcare standards. Surveys [15, 38, 9] highlight the potential and challenges of LLMs in healthcare, noting limited practical use. Our work contributes to advancing the practical deployment of LLMs for testing CRN software systems.
In the past, we developed a model-based framework to model the medical rules for GURI [43] with the Unified Modeling Language (UML) and Object Constraint Language (OCL).
Next, we developed an impact analysis approach [44] to study the impact of rule changes in response to GURI’s evolution (e.g., changes in regulations).
In addition, we developed a search-based approach to automatically refactor rules [31] to enhance rule understandability and maintainability.
However, these approaches do not help find problems in the modeled rules.
More recently, four studies advanced the testing at the CRN, mainly focusing on the GURI subsystem [27, 28, 24, 30].
@@ VII Related Work
Laaber et al. [28] address the practical complexities of testing an evolving system, offering insights applicable to healthcare registries.
Moreover, we studied EvoMaster [2] (an AI-based testing tool) to assess its effectiveness regarding code coverage, error detection, and domain-specific rule coverage at the CRN [27].
We extended EvoMaster and proposed EvoClass [24] which uses a ML classifier to significantly reduce cost without compromising effectiveness.
Lastly, EvoCLINICAL [30], a cyber-cyber digital twin for GURI, was proposed, which with fine-tuning through active transfer learning, can deal with system evolution.
In contrast to these works, this paper addresses the test oracle problem associated with rule execution results by employing differential testing.
Moreover, we tackle the challenge of generating electronic health records (EHRs) by leveraging the capabilities of LLMs.
VIII Conclusions
@@
Acknowledgments
The Norwegian Ministry of Education and Research supports Erblin Isaku’s PhD work reported in this paper. The research presented is also partly supported by The Research Council of Norway (RCN) under the project 309642 and has benefited from the Experimental Infrastructure for Exploration of Exascale Computing (eX3), which is supported by the RCN project 270053.